OpenStudy (anonymous):

What value of x makes the equation true? 2/8 = 12/?

OpenStudy (anonymous):

x=48 2/8 and 12/48

OpenStudy (anonymous):

2/8 = 12/x 2x = 96 x = 48

OpenStudy (anonymous):

The ? is supposed to be x right?

OpenStudy (anonymous):

or 2/8 = 12/48

OpenStudy (anonymous):

Thanks everyone :) :)

OpenStudy (anonymous):

welcome